Grade 9Trigonometry

solve (x+1) cot sq 30°=2cos sq 60°+3/4 sec sq 45°+4sin sq 30°

Anuradha Hela
8 Years agoGrade 9

1 Answer

Arun
8 Years ago
Dear student
remember cot 30 = sqrt (3)
cos 60 = ½
sec 45  = sqrt (2)
sin 30 = ½
now
(x+1)* 3 = 2* (1/4) + (3/4)* 2 + 4 * (1/4)
3 x+ 3 = ½ + 3/2 + 1
3 x = 3 – 3
3 x = 0
x = 0
Regards
Arun (askIITians forum expert)